# The Efficient Tail Hypothesis: An Extreme Value Perspective on Market Efficiency

### v1

@@ 5.4 Exploiting the Efficient Tail Hypothesis analysis
To motivate the practical application of the DTD measure, we construct an artificial dynamic portfolio based on the significant tail-inefficient pairs found in Section 5.3. For each pair satisfying $|\hat{\lambda}_{n}^{2}(X_{i},Y_{j})|\gg 0$, we suppose that we enter into the market for asset $Y_{j}$ (long position if $\hat{\lambda}_{n}^{2}(X_{i},Y_{j})>0$, otherwise short position) when we observe an extreme value for asset $X_{i}$ (given it exceeds its 99.5% historical quantile for positive extremes and falls below its 0.5% quantile when considering negative extremes). We hold the asset for one period and then exit the market by closing the position. With the trading cost ignored, we backtest this strategy on the out-of-sample data from 2023-06-01 to 2023-07-31. The cumulative Profit and Loss for each of the 126 significant pairs is given in Figure 5, alongside the overall Profit and Loss for the portfolio. We show a positive Profit and Loss throughout, suggesting that a trading strategy based on tail inefficiencies has the potential to make excess returns.
Figure 5: Profit and Loss of a dynamic portfolio consisting of the 126 tail-inefficient pairs. Upper panel shows the Profit and Loss for each pair and the lower panel shows the running cumulative Profit and Loss for the portfolio. The $y$-axis is the cumulative Profit and Loss (%) and the $x$-axis is the time index for the two-month test period.
@@ 6 Conclusion
In our data application, we investigate the market-wide tail efficiency of China’s futures market. Our empirical results suggest that the Efficient Tail Hypothesis (ETH) for China’s futures market is rejected under a significance level of $\alpha^{*}=0.01$. A deeper analysis of those tail-inefficient assets shows that negative extreme depreciation is more influential than positive extreme events. Moreover, we find evidence to support the momentum effect in extremes, which suggests that the market tends to be under-active in extreme situations. To capitalize on our findings, we construct an artificial dynamic portfolio consisting of those exceptional assets. The strategy is back-tested on out-of-sample data and shows a significant overall excess return. The results suggest that the analysis of directional tail dependence is useful for risk management and for constructing trading strategies in the financial market.
Several directions can be explored in future research. One avenue is to extend the concept of directional tail dependence to higher dimensions. Additionally, current causal inference methods in extremes can be expanded to consider both upper and lower tails, using our new directional tail dependence measure.
@@ Appendix A Proofs
Condition 1 follows by Lemma 6.1 of Resnick, (2007), which is a criterion for vague convergence to measures; It states that to show $\lim_{n\to\infty}n\mathbb{P}\{b_{n}^{-1}|\bm{X}|\in\cdot\}=\nu_{|\bm{X}|}(\cdot)$ on ${\mathbb{E}^{p}_{+}=\mathbb{R}_{+}\setminus\{\bm{0}\}}$, it is enough to show that $\lim_{n\to\infty}n\mathbb{P}\{b_{n}^{-1}|\bm{X}|\in[\bm{0},\bm{z}]^{c}\}=\nu_{|\bm{X}|}([\bm{0},\bm{z}]^{c})$ for all $\bm{z}:=(z_{1},\dots,z_{p})^{\top}\in[{0},\infty)^{p}\setminus\{\bm{0}\}$. We show that $\nu_{|\bm{X}|}(\cdot)$ can be expressed in terms of $\nu_{\bm{X}^{*}}(\cdot),$ which is the limit measure for $\bm{X}^{*}$. Consider
$\mathbb{P}\{b_{n}^{-1}|\bm{X}|$ $\in[\bm{0},\bm{z}]^{c}\}=\mathbb{P}\left\{\bm{X}\in\mathbb{R}^{p}:\bigvee_{i=1}^{p}\left[b_{n}^{-1}\left|X_{i}\right|>z_{i}\right]\right\}$
=$=\mathbb{P}\left\{\bm{X}^{*}\in\mathbb{R}^{p}:\bigvee_{i=1}^{p}\left[b_{n}^{-1}\left|\max\{(c_{i}^{+})^{-1/\alpha}X^{*}_{i},0\}+\min\{(c_{i}^{-})^{-1/\alpha}X_{i}^{*},0\}\right|>z_{i}\right]\right\}$
=$=\mathbb{P}\left\{\bm{X}^{*}\in\mathbb{R}^{p}:\bigvee_{i=1}^{p}\left[\left(b_{n}^{-1}(c_{i}^{+})^{-1/\alpha}X^{*}_{i}>z_{i}\right)\lor\left(b_{n}^{-1}(c_{i}^{-})^{-1/\alpha}X_{i}^{*}<-z_{i}\right)\right]\right\}$
=$=\mathbb{P}\left\{\bm{X}^{*}\in\mathbb{R}^{p}:\bigvee_{i=1}^{p}\left[\left(b_{n}^{-1}X^{*}_{i}>z_{i}(c^{+}_{i})^{1/\alpha}\right)\lor\left(b_{n}^{-1}X^{*}_{i}<-z_{i}(c_{i}^{-})^{1/\alpha}\right)\right]\right\}.$
@@
$\lim_{n\to\infty}n\mathbb{P}\{b_{n}^{-1}|\bm{X}|\in[\bm{0},\bm{z}]^{c}\}=\lim_{n\to\infty}n\mathbb{P}\left\{\bm{X}\in\mathbb{R}^{p}:\bigvee_{i=1}^{p}\left[b_{n}^{-1}\left|X_{i}\right|>z_{i}\right]\right\}$
$=$ $\lim_{n\to\infty}n\mathbb{P}\left\{\bm{X}^{*}\in\mathbb{R}^{p}:\bigvee_{i=1}^{p}\left[\left(b_{n}^{-1}X^{*}_{i}>z_{i}(c^{+}_{i})^{1/\alpha}\right)\lor\left(b_{n}^{-1}X^{*}_{i}<-z_{i}(c_{i}^{-})^{1/\alpha}\right)\right]\right\}$
$=$ $\nu_{\bm{X}^{*}}(\{\bm{x}\in\mathbb{E}^{p}:\bigvee_{i=1}^{p}x_{i}\in\{[-\infty,-z_{i}(c_{i}^{-})^{1/\alpha})\cup(z_{i}(c_{i}^{+})^{1/\alpha},\infty]\}\})=:\nu_{|\bm{X}|}([0,\bm{z}]^{c})$
@@
For Condition ii), we show that $\lim_{n\rightarrow\infty}\frac{n\mathbb{P}\{b_{n}^{-1}\bm{s}\odot\bm{X}\in[\bm{0},\bm{z}]^{c}\}}{n\mathbb{P}\{b_{n}^{-1}|\bm{X}|\in[\bm{0},\bm{z}]^{c}\}}\in(0,1)$ for all $\bm{z}\in[0,\infty)^{p}\setminus\{\bm{0}\}$ and any $\bm{s}\in\{-1,1\}^{p}$. Consider the numerator; we first show that this is strictly greater than zero. We have that $\lim_{n\rightarrow\infty}n\mathbb{P}\{\bm{s}\odot\bm{X}\in b_{n}[\bm{0},\bm{z}]^{c}\}\geq\lim_{n\rightarrow\infty}\mathbb{P}\{s_{1}b_{n}^{-1}X_{1}>z_{1}\}$, as the region $\{s_{1}b_{n}^{-1}X_{1}>z_{1}\}$ is a subspace of the region $\{\bm{s}\odot\bm{X}\in b_{n}[\bm{0},\bm{z}]^{c}\}$, and
$\lim_{n\rightarrow\infty}\mathbb{P}\{s_{1}b_{n}^{-1}X_{1}>z_{1}\}$ =$=\begin{cases}\lim_{n\rightarrow\infty}n\mathbb{P}\{b_{n}^{-1}X^{*}_{1}>z_{1}(c^{+}_{1})^{1/\alpha}\},&\text{if }s_{1}=1,\\ \lim_{n\rightarrow\infty}n\mathbb{P}\{b_{n}^{-1}X^{*}_{1}<-z_{1}(c^{-}_{1})^{1/\alpha}\},&\text{if }s_{1}=-1,\end{cases}$
@@
where the last line follows due to the homogeneity property of the limit measure. A similar argument can be used to show that $\nu_{\bm{X}}(\{\bm{x}\in\mathbb{E}^{p}:x_{i}<-1\})=1$ for any $i=1,\dots,p$. Thus, all three conditions are satisfied and this completes the proof.
@@
We first show that $\bar{\bm{X}}:=(\bar{X}_{1},\dots,\bar{X}_{2p})^{\top}\in{\rm RV}^{2p}_{+}(2)$ with limit measure $\nu_{\bar{\bm{X}}}(\cdot)$, which we express in terms of $\nu_{{\bm{X}}}(\cdot)$. By Lemma 6.1 of Resnick, (2007), we have, for all $\bm{\bar{z}}:=((\bm{z^{+}})^{\top};(\bm{z^{-}})^{\top})^{\top}\in[0,\infty)^{2p}\setminus\{\bm{0}\}$, that
$\lim_{n\to\infty}n\mathbb{P}\{b_{n}^{-1}\bar{\bm{X}}\in[\bm{0},\bm{\bar{z}}]^{c}\}$ $=\lim_{n\to\infty}n\mathbb{P}\{\bigvee_{i=1}^{p}[X^{+}_{i}>z^{+}_{i}\lor X^{-}_{i}>z^{-}_{i}]\}$
$=\lim_{n\to\infty}n\mathbb{P}\{\bigvee_{i=1}^{p}[\max\{X_{i},0\}>z^{+}_{i}\lor-\min\{X_{i},0\}>z^{-}_{i}]\}$
$=\lim_{n\to\infty}n\mathbb{P}\{\bigvee_{i=1}^{p}[X_{i}\in(-\infty,-z^{-}_{i})\cup(z^{+}_{i},\infty)\}$
$=\nu_{\bm{X}}(\{\bigvee_{i=1}^{p}[(-\infty,-z^{-}_{i})\cup(z^{+}_{i},\infty)]\})=:\nu_{\bar{\bm{X}}}([\bm{0},\bm{\bar{z}}]^{c}).$
Thus, $\bar{\bm{X}}\in{\rm RV}^{2p}_{+}(2)$ with limit measure $\nu_{\bar{\bm{X}}}([\bm{0},\bm{\bar{z}}]^{c})=\nu_{\bm{X}}(\{\bigvee_{i=1}^{p}[(-\infty,-z^{-}_{i})\cup(z^{+}_{i},\infty)]\})$.
@@
$\nu_{\bar{\bm{X}}}(\{\bm{x}\in\mathbb{E}^{2p}_{+}:x_{i}>1\})$ $=\lim_{n\rightarrow\infty}\mathbb{P}\{b_{n}^{-1}\bar{X}_{i}>1\}=\lim_{n\rightarrow\infty}\mathbb{P}\{b_{n}^{-1}\max\{X_{i},0\}>1\}$
@@
$=\lim_{n\rightarrow\infty}\mathbb{P}\{b_{n}^{-1}X_{i}>1\}=\nu_{\bm{X}}(\{\bm{x}\in\mathbb{E}^{p}:x_{i}>1\})=1,$
@@
$\nu_{\bar{\bm{X}}}(\{\bm{x}\in\mathbb{E}^{2p}_{+}:x_{i+p}>1\})$ $=\lim_{n\rightarrow\infty}\mathbb{P}\{-b_{n}^{-1}\bar{X}_{i}>1\}=\lim_{n\rightarrow\infty}\mathbb{P}\{-b_{n}^{-1}\min\{X_{i},0\}>1\}$
$=\lim_{n\rightarrow\infty}\mathbb{P}\{b_{n}^{-1}X_{i}<-1\}=\nu_{\bm{X}}(\{\bm{x}\in\mathbb{E}^{p}:x_{i}<-1\})=1.$
@@
A.4 Proof of directional tail dependence properties
We first show that $\lambda(X,Y)\in[-1,1]$, where $\lambda(X,Y)=-1$ and $\lambda(X,Y)=1$ are achieved when $\lim_{a\rightarrow\infty}\mathbb{P}[X>a\mid Y<-a]=1$ and $\lim_{a\rightarrow\infty}\mathbb{P}[X>a\mid Y>a]=1$, respectively. By Proposition 3.4, we have that $\lambda(X,Y)=2\{\sigma(X^{+},Y^{+})-\sigma(X^{+},Y^{-})\}$. As $\sigma(X^{+},Y^{+})\in[0,{1\over 2}]$ (and similarly for $\sigma(X^{+},Y^{-})$), we have that $\lambda(X,Y)\in[-1,1]$. The boundary case $\lambda(X,Y)=1$ implies $\sigma(X^{+},Y^{+})=0.5$ and $\theta_{X^{+},Y^{+}}=0$ which further implies $\sigma(X^{+},Y^{-})=0$ as $\theta_{X^{+},Y^{-}}\in[\frac{\pi}{2}-\theta_{X^{+},Y^{+}},\frac{\pi}{2}]$ by Proposition 3.3. The case $\sigma(X^{+},Y^{+})=0.5$ is achieved when
@@ A.4 Proof of directional tail dependence properties
The third line follows as $X^{+}=X$ and $Y^{+}=Y$ for any $(X,Y)^{\top}$ such that ${{(X,Y)^{\top}\over{||(X,Y)^{\top}||_{2}}}\in\mathring{\mathbb{S}}^{1}_{+}}$. A similar argument can be used to show that, for $\Lambda\subseteq\mathring{\mathbb{S}}^{1}_{+}$, we have $H_{(X,Y)}(\Lambda^{\prime})=H_{X^{+},Y^{-}}(\Lambda)$.
@@
The Proof of Proposition 3.5 follows the same strategy as in Larsson and Resnick, (2012). The major difference comes from the estimator’s definition of an “exceedance”. In our case, we use $\mathbbm{1}\{r_{i}\geq r_{0}\}$ where $r_{0}=a(n/k)$ (same setting as used in Resnick, (2004)), while in Larsson and Resnick, (2012), they use $\mathbbm{1}\{r_{i}\geq r_{(k)}\}$ where $r_{(k)}$ is the $k$-th order statistic of $\left\{r_{i}\right\}_{i=1}^{n}$.
See 3.5

### v2

@@ 5.4 Exploiting the Efficient Tail Hypothesis analysis
To motivate the practical application of the DTD measure, we construct an artificial dynamic portfolio based on the significant tail-inefficient pairs found in Section 5.3. Each inefficient pair falls into one of the following four cases: i) $\hat{\lambda}_{n}^{2}(Z_{i,t},Z_{j,t+1})>0$: an extremal appreciation event for asset $i$ is associated with an extremal appreciation event for asset $j$ in the next timestep; ii) $\hat{\lambda}_{n}^{2}(Z_{i,t},Z_{j,t+1})<0$: an extremal appreciation event for asset $i$ is associated with an extremal depreciation event for asset $j$ in the next timestep; iii) $\hat{\lambda}_{n}^{2}(-Z_{i,t},Z_{j,t+1})>0$: an extremal depreciation event for asset $i$ is associated with an extremal appreciation event for asset $j$ in the next timestep; iv) $\hat{\lambda}_{n}^{2}(-Z_{i,t},Z_{j,t+1})<0$: an extremal depreciation event for asset $i$ is associated with an extremal depreciation event for asset $j$ in the next timestep.
For each pair satisfying $|\hat{\lambda}_{n}^{2}(X_{i},Y_{j})|>0$, we suppose that we enter into the market for asset $Y_{j}$ (long position if $\hat{\lambda}_{n}^{2}(X_{i},Y_{j})>0$, otherwise short position) when we observe an extreme value for asset $X_{i}$ (given it exceeds its 99.5% historical quantile for positive extremes and falls below its 0.5% quantile when considering negative extremes). We hold the asset for one period and then exit the market by closing the position. With the trading cost ignored, we backtest this strategy on the out-of-sample data from 2023-06-01 to 2023-07-31. The cumulative Profit and Loss for each of the 60 significant pairs is given in Figure 5, alongside the overall Profit and Loss for the portfolio. We also study the case when set significance level to $0.05$ under which we find 294 significant pairs. Both results show a positive Profit and Loss throughout, suggesting that a trading strategy based on tail inefficiencies has the potential to make excess returns. This result is consistent with the bootstrap experiment, which is deferred to Supplementary Material, Section D.4.
As stated by the Efficient Market Hypothesis, past information cannot be used to predict future price movements. However, our backtesting on out-of-sample data suggests that our trading strategy using tail-inefficient assets is profitable (under the ideal situation where the trading cost is not considered), which gives evidence against the Efficient Market Hypothesis. Therefore, the Efficient Tail Hypothesis, which is testable, can be viewed as an analogue of the Efficient Market Hypothesis with a stronger focus on extremal events.
Figure 5: Profit and Loss of a dynamic portfolio consisting of the 60 or 294 tail-inefficient pairs under significance level $\alpha^{*}=0.01$ (left) or $\alpha^{*}=0.05$ (right), respectively. Upper panels show the Profit and Loss for each pair and the lower panels show the running cumulative Profit and Loss for the portfolio. The $y$-axis is the cumulative Profit and Loss (%) and the $x$-axis is the time index for the two-month test period.
@@ 6 Conclusion
In our data application, we investigate the market-wide tail efficiency of China’s futures market. Our empirical results suggest that the Efficient Tail Hypothesis (ETH) for China’s futures market is rejected under a significance level of $\alpha^{*}=0.01$. A deeper analysis of those tail-inefficient assets shows that negative extreme depreciation is more influential than positive extreme events. Moreover, we find evidence to support the momentum effect in extremes, which suggests that the market tends to be under-active in extreme situations. To capitalize on our findings on momentum effect in extremes, we construct an artificial dynamic portfolio consisting of those exceptional assets. The strategy is back-tested on out-of-sample data and shows a significant overall excess return. Those results suggest that most one-step forward extremal dependencies are directionally tail symmetric, while significant deviations indicate potential trading opportunities. Therefore, the analysis of directional tail dependence is useful for risk management and for constructing trading strategies in the financial market.
Several directions can be explored in future research. One avenue is to extend the concept of directional tail dependence to higher dimensions. Additionally, current causal inference methods in extremes can be expanded to consider both upper and lower tails, using our new directional tail dependence measure. Last but not least, further investigation into simulation methods for time series data with specific directional tail dependence structures could enable the generation of time series data with varying degrees of directional asymmetry in financial applications.
@@
Supplementary Material for “The Efficient Tail Hypothesis: An Extreme Value Perspective on Market Efficiency”
Introduction
This Supplementary Material provides additional details and results for the main paper. Section A contains the proofs. Section B describes the dataset used in the Application section. Section C provides additional simulation results for the scenario with smaller sample sizes and lower exceedance thresholds. Finally, Section D presents further results for the empirical application.
@@ Appendix A Proofs
[…]
$\mathbb{P}\{b_{n}^{-1}|\bm{X}|$ $\in[\bm{0},\bm{z}]^{c}\}=\mathbb{P}\left\{\bm{X}\in\mathbb{R}^{p}:{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\bigcup}_{i=1}^{p}\left[b_{n}^{-1}\left|X_{i}\right|>z_{i}\right]\right\}$
=$=\mathbb{P}\left\{\bm{X}^{*}\in\mathbb{R}^{p}:{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\bigcup}_{i=1}^{p}\left[b_{n}^{-1}\left|\max\{(c_{i}^{+})^{-1/\alpha}X^{*}_{i},0\}+\min\{(c_{i}^{-})^{-1/\alpha}X_{i}^{*},0\}\right|>z_{i}\right]\right\}$
=$=\mathbb{P}\left\{\bm{X}^{*}\in\mathbb{R}^{p}:{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\bigcup}_{i=1}^{p}\left[\left(b_{n}^{-1}(c_{i}^{+})^{-1/\alpha}X^{*}_{i}>z_{i}\right){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\cup}\left(b_{n}^{-1}(c_{i}^{-})^{-1/\alpha}X_{i}^{*}<-z_{i}\right)\right]\right\}$
=$=\mathbb{P}\left\{\bm{X}^{*}\in\mathbb{R}^{p}:{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\bigcup}_{i=1}^{p}\left[\left(b_{n}^{-1}X^{*}_{i}>z_{i}(c^{+}_{i})^{1/\alpha}\right){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\cup}\left(b_{n}^{-1}X^{*}_{i}<-z_{i}(c_{i}^{-})^{1/\alpha}\right)\right]\right\}.$
@@
$\lim_{n\to\infty}n\mathbb{P}\{b_{n}^{-1}|\bm{X}|\in[\bm{0},\bm{z}]^{c}\}=\lim_{n\to\infty}n\mathbb{P}\left\{\bm{X}\in\mathbb{R}^{p}:{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\bigcup}_{i=1}^{p}\left[b_{n}^{-1}\left|X_{i}\right|>z_{i}\right]\right\}$
$=$ $\lim_{n\to\infty}n\mathbb{P}\left\{\bm{X}^{*}\in\mathbb{R}^{p}:{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\bigcup}_{i=1}^{p}\left[\left(b_{n}^{-1}X^{*}_{i}>z_{i}(c^{+}_{i})^{1/\alpha}\right){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\cup}\left(b_{n}^{-1}X^{*}_{i}<-z_{i}(c_{i}^{-})^{1/\alpha}\right)\right]\right\}$
$=$ $\nu_{\bm{X}^{*}}(\{\bm{x}\in\mathbb{E}^{p}:{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\bigcup}_{i=1}^{p}x_{i}\in\{[-\infty,-z_{i}(c_{i}^{-})^{1/\alpha})\cup(z_{i}(c_{i}^{+})^{1/\alpha},\infty]\}\})=:\nu_{|\bm{X}|}([0,\bm{z}]^{c})$
@@
For Condition ii), we show that $\lim_{n\rightarrow\infty}\frac{n\mathbb{P}\{b_{n}^{-1}\bm{s}\odot\bm{X}\in[\bm{0},\bm{z}]^{c}\}}{n\mathbb{P}\{b_{n}^{-1}|\bm{X}|\in[\bm{0},\bm{z}]^{c}\}}\in(0,1)$ for all $\bm{z}\in[0,\infty)^{p}\setminus\{\bm{0}\}$ and any $\bm{s}\in\{-1,1\}^{p}$. Consider the numerator; we first show that this is strictly greater than zero. We have that $\lim_{n\rightarrow\infty}n\mathbb{P}\{\bm{s}\odot\bm{X}\in b_{n}[\bm{0},\bm{z}]^{c}\}\geq\lim_{n\rightarrow\infty}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}n}\mathbb{P}\{s_{1}b_{n}^{-1}X_{1}>z_{1}\}$, as the region $\{s_{1}b_{n}^{-1}X_{1}>z_{1}\}$ is a subspace of the region $\{\bm{s}\odot\bm{X}\in b_{n}[\bm{0},\bm{z}]^{c}\}$, and
$\lim_{n\rightarrow\infty}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}n}\mathbb{P}\{s_{1}b_{n}^{-1}X_{1}>z_{1}\}$ =$=\begin{cases}\lim_{n\rightarrow\infty}n\mathbb{P}\{b_{n}^{-1}X^{*}_{1}>z_{1}(c^{+}_{1})^{1/\alpha}\},&\text{if }s_{1}=1,\\ \lim_{n\rightarrow\infty}n\mathbb{P}\{b_{n}^{-1}X^{*}_{1}<-z_{1}(c^{-}_{1})^{1/\alpha}\},&\text{if }s_{1}=-1,\end{cases}$
@@
where the last line follows due to the homogeneity property of the limit measure. A similar argument can be used to show that $\nu_{\bm{X}}(\{\bm{x}\in\mathbb{E}^{p}:x_{i}<-1\})=1$ for any $i=1,\dots,p$. Thus, all three conditions are satisfied and this completes the proof. ∎
@@
[…]
$\lim_{n\to\infty}n\mathbb{P}\{b_{n}^{-1}\bar{\bm{X}}\in[\bm{0},\bm{\bar{z}}]^{c}\}$ $=\lim_{n\to\infty}n\mathbb{P}\{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\bigcup}_{i=1}^{p}[X^{+}_{i}>z^{+}_{i}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\cup}X^{-}_{i}>z^{-}_{i}]\}$
$=\lim_{n\to\infty}n\mathbb{P}\{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\bigcup}_{i=1}^{p}[\max\{X_{i},0\}>z^{+}_{i}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\cup}-\min\{X_{i},0\}>z^{-}_{i}]\}$
$=\lim_{n\to\infty}n\mathbb{P}\{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\bigcup}_{i=1}^{p}[X_{i}\in(-\infty,-z^{-}_{i})\cup(z^{+}_{i},\infty)\}$
$=\nu_{\bm{X}}(\{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\bigcup}_{i=1}^{p}[(-\infty,-z^{-}_{i})\cup(z^{+}_{i},\infty)]\})=:\nu_{\bar{\bm{X}}}([\bm{0},\bm{\bar{z}}]^{c}).$
Thus, $\bar{\bm{X}}\in{\rm RV}^{2p}_{+}(2)$ with limit measure $\nu_{\bar{\bm{X}}}([\bm{0},\bm{\bar{z}}]^{c})=\nu_{\bm{X}}(\{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\bigcup}_{i=1}^{p}[(-\infty,-z^{-}_{i})\cup(z^{+}_{i},\infty)]\})$.
@@
$\nu_{\bar{\bm{X}}}(\{\bm{x}\in\mathbb{E}^{2p}_{+}:x_{i}>1\})$ $=\lim_{n\rightarrow\infty}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}n}\mathbb{P}\{b_{n}^{-1}\bar{X}_{i}>1\}=\lim_{n\rightarrow\infty}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}n}\mathbb{P}\{b_{n}^{-1}\max\{X_{i},0\}>1\}$
@@
$=\lim_{n\rightarrow\infty}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}n}\mathbb{P}\{b_{n}^{-1}X_{i}>1\}=\nu_{\bm{X}}(\{\bm{x}\in\mathbb{E}^{p}:x_{i}>1\})=1,$
@@
$\nu_{\bar{\bm{X}}}(\{\bm{x}\in\mathbb{E}^{2p}_{+}:x_{i+p}>1\})$ $=\lim_{n\rightarrow\infty}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}n}\mathbb{P}\{-b_{n}^{-1}\bar{X}_{i}>1\}=\lim_{n\rightarrow\infty}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}n}\mathbb{P}\{-b_{n}^{-1}\min\{X_{i},0\}>1\}$
$=\lim_{n\rightarrow\infty}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}n}\mathbb{P}\{b_{n}^{-1}X_{i}<-1\}=\nu_{\bm{X}}(\{\bm{x}\in\mathbb{E}^{p}:x_{i}<-1\})=1.$
@@
A.4 Proof of Proposition 3.4
See 3.4 We first show that $\lambda(X,Y)\in[-1,1]$, where $\lambda(X,Y)=-1$ and $\lambda(X,Y)=1$ are achieved when $\lim_{a\rightarrow\infty}\mathbb{P}[X>a\mid Y<-a]=1$ and $\lim_{a\rightarrow\infty}\mathbb{P}[X>a\mid Y>a]=1$, respectively. By Proposition 3.5, we have that $\lambda(X,Y)=2\{\sigma(X^{+},Y^{+})-\sigma(X^{+},Y^{-})\}$. As $\sigma(X^{+},Y^{+})\in[0,{1\over 2}]$ (and similarly for $\sigma(X^{+},Y^{-})$), we have that $\lambda(X,Y)\in[-1,1]$. The boundary case $\lambda(X,Y)=1$ implies $\sigma(X^{+},Y^{+})=0.5$ and $\theta_{X^{+},Y^{+}}=0$ which further implies $\sigma(X^{+},Y^{-})=0$ as $\theta_{X^{+},Y^{-}}\in[\frac{\pi}{2}-\theta_{X^{+},Y^{+}},\frac{\pi}{2}]$ by Proposition 3.3. The case $\sigma(X^{+},Y^{+})=0.5$ is achieved when
@@
The third line follows as $X^{+}=X$ and $Y^{+}=Y$ for any $(X,Y)^{\top}$ such that ${{(X,Y)^{\top}\over{||(X,Y)^{\top}||_{2}}}\in\mathring{\mathbb{S}}^{1}_{+}}$. A similar argument can be used to show that, for $\Lambda\subseteq\mathring{\mathbb{S}}^{1}_{+}$, we have $H_{(X,Y)}(\Lambda^{\prime})=H_{X^{+},Y^{-}}(\Lambda)$. ∎
@@
[…]
See 3.6